OpenStudy (anonymous):

please please help me on this one? which of the following represents the radical in simpleast form? √(b^14) a.b^7 b. √(b^7) c. b^7 √(b)

OpenStudy (anonymous):

b^(m+n) = (b^m)^n. If you have sqrt(x^2) , then that is equal to x.

OpenStudy (anonymous):

argh, I mean b^(m*n) = (b^m)^n ... and 14 = 7*2.

OpenStudy (anonymous):

so which one would it be?

OpenStudy (anonymous):

If you are enrolled in the class, then you should rightly be able to tell me. I've given you a pretty good hint!

OpenStudy (anonymous):

You need to think about it.